OpenStudy (anonymous):

Calculate the circumference of a circle with radius 5.

Parth (parthkohli):

\(C = 2\pi r \Longrightarrow 2 \times 3.14 \times 5\)

Parth (parthkohli):

Welcome to OpenStudy!

OpenStudy (anonymous):

thnx